# OBsmith: LLM-powered JavaScript Obfuscator Testing

### v1

@@ Abstract.
JavaScript obfuscators are widely deployed to protect intellectual property and resist reverse engineering, yet their correctness has been largely overlooked compared to performance and resilience. Existing evaluations typically measure resistance to deobfuscation, leaving the critical question of whether obfuscators preserve program semantics unanswered. Incorrect transformations can silently alter functionality, compromise reliability, and erode security—undermining the very purpose of obfuscation. To address this gap, we present OBsmith, a novel framework to systematically test JavaScript obfuscators using large language models (LLMs). OBsmith leverages LLMs to generate program sketches—abstract templates capturing diverse language constructs, idioms, and corner cases—which are instantiated into executable programs and subjected to obfuscation under different configurations. Besides LLM-powered sketching, OBsmith also employs a second source: automatic extraction of skeletons from real programs. This extraction path enables more focused testing of project-specific features and lets developers inject domain knowledge into the resulting test cases. OBsmith uses two techniques to derive test oracles: (i) differential testing, which treats the original program as ground truth and instruments control- and data-flow to detect semantic divergences, failures, or crashes; and (ii) metamorphic testing, which applies semantics-preserving transformations to the original program and checks if obfuscation violates expected behavior.
We evaluate OBsmith on two widely used obfuscators, Obfuscator.IO and JS-Confuser, generating 600 sketches using six popular LLMs. OBsmith fills these sketches and generates over 3,000 candidate programs and obfuscates them across seven obfuscation configurations. OBsmith uncovers 11 previously unknown correctness bugs. Under an equal program budget, five general purpose state-of-the-art JavaScript fuzzers (FuzzJIT, Jsfunfuzz, Superion, DIE, Fuzzilli) failed to detect these issues, highlighting OBsmith’s complementary focus on obfuscation-induced misbehavior. An ablation shows that all components except our generic MRs contribute to at least one bug class; the negative MR result suggests the need for obfuscator-specific metamorphic relations. Our results also seed discussion on how to balance obfuscation presets and performance cost. We envision OBsmith as an important step towards automated testing and quality assurance of obfuscators and other semantic-preserving toolchains.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Formal software verificationccs: Software and its engineering Correctnessccs: Software and its engineering Completenessccs: Software and its engineering Software verificationccs: Computing methodologies Artificial intelligenceccs: General and reference Metricsccs: General and reference Evaluation
@@ 1. Introduction
By design, an effective obfuscator is expected to complicate the original source code without altering its intended functionality. The primary goal is to make the code substantially more difficult for reverse engineers and automated analysis tools to understand, thus safeguarding proprietary or sensitive information. Ensuring functional equivalence between obfuscated and original code is critical, as obfuscation may introduce unintended behaviors, potentially undermining software reliability or security. There is still a lack of research for evaluating Obfuscators on pure JavaScript code. Previous research on JavaScript obfuscation largely focuses on evaluating the effectiveness of obfuscation techniques or detecting if code is obfuscated. However, comprehensive testing frameworks that validate both functional correctness and runtime performance impacts of JavaScript obfuscators remain underexplored. Recent research by Skolka et al.(Skolka et al., 2019) offers a comprehensive evaluation of JavaScript minifiers and obfuscators, with particular emphasis on their robustness when applied to client-side scripts. The authors primarily relied on unit test to assess the effectiveness of various obfuscators. Their findings reveal that these tools often exhibit insufficient resilience. A core challenge of client-side script is that client-side JavaScript is inherently dependent on complex factors such as user interactions, asynchronous events, and integration with real-world browser APIs (e.g., DOM, cookies, localStorage). These dependencies are difficult to replicate or comprehensively test using automated methods. As a result, automated source code transformations—such as minification and obfuscation—may inadvertently introduce bugs in client side. Moreover, prior studies have omitted the evaluation of some of the most widely adopted obfuscators, such as Obfuscator.IO(obi, 2024) and JS-Confuser(jsc, 2022), as indicated by a recent Google survey(Jiang et al., 2025). This omission constitutes another important research gap, particularly given the pervasive use of these tools in the software industry. To address this limitation, the present study aims to instrument and evaluate pure JavaScript code that is executable in both Node.js and browser environments. Addressing this research gap is crucial to ensuring that JavaScript obfuscators can reliably achieve their intended security and functionality objectives without inadvertently compromising software behavior.
@@
General purpose compiler testing approaches lack of a systematically-enhanced test oracle to capture the JavaScript obfuscator bugs. Off-the-shelf compiler fuzzers assume closed-world, deterministic pipelines whose goal is transparent semantics preservation; they rely on crash oracles or n-version agreement (e.g., GCC vs. Clang). JavaScript obfuscators instead preserve while disguising semantics, inject seed-dependent randomness and self-defense checks, and interact with host features that general fuzzers neither generate nor control. Bugs often surface in JS-specific corners—eval function, dynamic code loading, scoping/hoisting, prototype mutations, getters/setters, proxies—combined with obfuscation patterns like control-flow flattening or lazy string decoding. These properties defeat differential comparisons across tools and make crashes a poor proxy for silent miscompilations. We take the original program as the ground-truth oracle and check semantic equivalence—including return values, exceptions, scheduling-sensitive effects, and observable host-API interactions. OBsmith generates JS-centric, event-aware sketches, controls nondeterminism via seed management, executes in sandboxed Node and browser-like environments, and is configuration-aware to sweep obfuscator options. Lightweight instrumentation records variable states and control-flow to localize divergences. This semantics-, host-, and transformation-aware design exposes obfuscation-specific bugs that general compiler fuzzers systematically miss. Obfuscators differ from compilers because they preserve but disguise semantics. They often use tricks (control-flow flattening, dead code) that do not appear in compiler pipelines. Hence, domain-specific sketching/testing is required.
@@
We propose OBsmith, a novel testing framework specifically designed for evaluating JavaScript obfuscators. OBsmith leverages sketching techniques to systematically generate diverse test programs and employs differential testing to identify discrepancies introduced by obfuscation processes. Sketching and differential testing methodologies have previously demonstrated effectiveness in compiler testing, enabling the detection of subtle semantic bugs and inconsistencies. Extensive literature on compiler testing underscores the value and efficacy of integrating domain-specific knowledge into test-case generation. Inspired by these insights, we adopt sketching to incorporate JavaScript-specific domain knowledge into our framework, thereby enhancing the capability of OBsmith to detect functional deviations and performance degradations caused by obfuscators.
@@
OBsmith uses LLM to generate sketches and uses Babel(bab, 2014) to implement the verification and differential testing mechanism. Trained on extensive textual corpora, LLMs inherently encapsulate rich domain-specific knowledge, making them particularly suitable for generating sketches in automated test-case generation. Leveraging their ability to understand syntactic structures and semantic constraints, LLMs can systematically produce representative code sketches that reflect realistic programming patterns and corner cases. Integrating LLM-based sketch generation into differential testing frameworks enhances test diversity and coverage with low resource, thus enabling the effective discovery of subtle functional inconsistencies and edge-case defects in software, such as those potentially introduced by JavaScript obfuscators.
@@
A critical component of OBsmith is its integrated program generator, which is engineered to rigorously explore different execution path and expose the semantic equivalence between the original source program and its obfuscated counterpart. This step is indispensable, as the intricate nature of obfuscation transformations carries an inherent risk of introducing functional regressions or altering program behavior in subtle ways. To mitigate this risk, OBsmith contains a sketch filling algorithm to generate a comprehensive set of inputs to test different execution paths. OBsmith uses a program enhancer to record global and local variables and record the program’s control flow. For each input, OBsmith asserts that the original and obfuscated program versions produce identical outputs and exhibit equivalent observable side-effects. This process provides a strong, empirical guarantee that the core logic and functionality of the application remain unaltered post-obfuscation, thereby enhancing the trustworthiness and practical deployability of our approach.
@@
Unlike the standard differential testing setup used in compiler validation – where multiple variants of a system are run on the same input and their outputs compared for inconsistencies – OBsmith  relies on a concrete ground truth. In particular, we treat the original, unobfuscated program as the definitive oracle for correct behavior. This direct comparison against a known-correct baseline allows for a precise and conclusive assessment of the correctness of the obfuscation transformations, ensuring that they have not introduced functional changes.
@@
We evaluate OBsmith on two widely used obfuscators, Obfuscator.IO and JS-Confuser, generating 600 sketches using six popular LLMs. OBsmith fills theese sketches and generates over 3,000 candidate programs and obfuscate them across seven obfuscation configurations. OBsmith uncovers 11 previously unknown correctness bugs. Under an equal program budget, five general purpose state-of-the-art JavaScript fuzzers (FuzzJIT, Jsfunfuzz, Superion, DIE, Fuzzilli) failed to rediscover these issues, highlighting OBsmith’s complementary focus on obfuscation-induced misbehavior. We also discuss obfuscation affect on file size, run time, and memory usage in discussion (§5).
@@
  • Novelty. This paper introduces OBsmith, the first LLM-powered framework for systematically testing JavaScript obfuscators. OBsmith introduces an LLM-driven, sketch-based generator (mixing LLM-created and automatically extracted sketches) plus a PL-aware oracle that combines differential testing of outputs/exceptions/termination with lightweight instrumentation and exploratory metamorphic tests.
  • Implementation. We implement OBsmith with 2 sketch source: (i) leverages multi-agent LLM system to automatically generate diverse sketches and (ii) use existing JavaScript programs automatically extract sketch. To solve oracle problem, OBsmith applies (i) differential testing and metamorphic testing to evaluate the correctness of obfuscation.
@@ 2. Sketch Example
In this paper, a sketch refers to a “program with holes,” where key components - such as expressions, variables, or literals - are left as placeholders to be filled in later. Sketches, also known as “skeletal programs” or “templates,” have been shown in previous work to be an effective mechanism for introducing domain knowledge into compiler testing(Zang et al., 2023; Zhang et al., 2017; Wang et al., 2018). By abstracting certain program elements, sketches allow for a flexible yet systematic exploration of diverse code scenarios that are effective in exposing compiler bugs. The left part of Figure 1 is an example of sketch with holes (e.g. NumberLiteral, NumberReference) and the middle is the corresponding filled code.
@@
JavaScript obfuscators take two inputs: the program under obfuscation and the obfuscation configuration. Once we obfuscate the concrete program with different obfuscators and different configurations, we got the multiple obfuscated programs. Then we conduct a differential testing on these programs to compare if they have the same functionality. Unlike traditional diff testing in compiler testing, we have ground truth (unobfuscated program). Here the concrete program serve as ground truth and we compare the running result to find bugs. Furthermore, OBsmith enhances programs by recording the program’s workflow and data flow, which effectively reflects the program functionality and enables us to easily implement differential testing. An example of found bug is shown in the right part of Figure 1. Here the variable x is reported as undefined during execution (ReferenceError: x is not defined). This error highlights a critical functionality bug introduced by JS-Confuser, demonstrating how sketch-based differential testing effectively detects discrepancies between original program and obfuscated program.
@@ 3. OBsmith Approach
In this section, we present OBsmith, an automated framework that leveraging LLMs to generate sketches and use skteches to test JavaScript obfuscators. OBsmith constructs a multi-step pipeline to test JavaScript obfuscators, its overall workflow is shown in Figure 2.
@@
Then LLM-generated and extracted sketches are used in program generation3.5) to get concrete JavaScript programs. In program generation stage, OBsmith (i) solves the sketch by filling all placeholders and generate concrete expressions in the sketch; and (ii) enhances the filled program to log the program’s control flow and data flow. The output of program generation stage is a set of enhanced candidate programs ready for obfuscation.
The last step is use candidate programs and call obfuscators to get multiple obfuscated variants. These variants with corresponding ground truth are input of the follow up testing stage. OBsmith contains two test oracles: differential testing3.6) and metamorphic testing3.7). OBsmith uses different JavaScript obfuscators with different obfuscation configurations to obfuscate candidate programs. The original candidate program and all obfuscated versions are executed within a standard JavaScript engine (Node.js). Unlike traditional differential testing in compiler testing which lacks of ground truth, OBsmith uses the original program as ground truth. OBsmith collects the execution outputs of all obfuscated programs and compare these outputs with the ground truth. If OBsmith finds any output inconsistencies, it reports the obfuscator as failed. How to define output inconsistencies is in §3.6.
@@ 3.1. Sketch Definition
Sketch, also known as ”template” or ”skeletal program”, has proved to be effective in compiler testing and in different programming languages (Zhang et al., 2017; Zang et al., 2023). Previous works define many Domain Specific Languages (DSLs) to represent different sketch syntax, but overall sketch represent a blueprint from which a multitude of concrete, executable programs can be systematically generated. An example of OBsmith’s sketch is shown in §2. OBsmith’s sketch is an abstract template for JavaScript programs, and we provide a powerful sketch syntax that supports abstractions for numbers, booleans, references, and expressions.
Sketch Syntax
This section provides a formal description of the sketch syntax, and its role within the program generation pipeline (filling and enhancement) that produces the final candidate programs used for differential testing of JavaScript Obfuscators. The primary goal of our sketching language is to enable the test program generation to be both diverse and controlled.
(1) Diversity: The syntax allows programs generation with varied control flows, data flows, and uses of language features. We can ensure the diversity by using same sketch with randomly fill placeholders. This can leads to different control flow and data flow.
(2) Control: Unlike purely random program generation, sketching allows the developers to define a specific program structure or to target a particular language feature, which is useful to inject their domain knowledge. This is essential for creating effective test cases that can isolate and identify bugs related to specific semantic constructs.
By design, our sketch is a syntactically valid JavaScript file where the placeholders are treated as if they were standard identifiers, literals, or expressions. This ensures that the sketch itself can be parsed by standard JavaScript tools (e.g. Babel), simplifying the program generation implementation.
Literal placeholders
are syntactic constructs that represent a location in the code where a new, randomly generated primitive value will be inserted by the filling script. OBsmith supports number and boolean literals now.
A numberLiteral represents a placeholder for a concrete numerical value. It can be float or integer (in JavaScript, the only data type is Number, no distinction between integers and floating-point numbers) and replaced with a randomly selected integer or floating-point number (e.g., 10, -42, 3.14159) in the following fill process. So it must be used in a context where a numerical value is syntactically valid.
A booleanLiteral represents a placeholder for a concrete boolean value. It is replaced with either true or false. Must be used in a boolean context, such as in conditional statements or variable assignments.
Reference placeholders
are used to create data flow dependencies in the generated program. The shim replaces these placeholders with references to existing variables of the corresponding type in the current scope. If no such variable exists in the current scope, the generator will throw an error to ensure semantic validity. OBsmith  supports both number and boolean references.
A numberReference represents a reference placeholder that points to an existing numeric variable in scope. It is replaced by the identifier name of a numeric variable. It is used to create a data dependency on a previously defined numeric variable.
A booleanReference represents a reference placeholder that points to an existing Boolean variable in scope. It is replaced by the identifier name of a Boolean variable. It is used to create a data dependency on a previously defined Boolean variable.
Expression placeholders
are used to represent a concrete logic, arithmetic or relation expression. The general expression placeholder is in the following form
expression_type(operand1, operand2, operator1, [operator2, ...])
For expression type,OBsmith supports arithmetic, relation, and logic. An arithmetic expression generates a binary arithmetic expression. The first two parameters represent the operands of the expression. The expression must resolve to a numeric value to be semantically valid at runtime. The following op1, [op2, …] can be one or more possible arithmetic operators. The generator randomly selects an operator from this set. A relation expression generates a binary relational expression that evaluates to a Boolean value. This is the primary construct for creating dynamic control flow. Operand 1 and operand 2 also represent the operands to be compared. They can be of any type. The following op1, [op2, …] can be one or more possible relational operators. The generator randomly selects one. A logic expression generates a binary logical expression. Operand 1 and operand 2 represent the operands of the logical operation. These operands should resolve to Boolean values. The following op1, [op2, …] can be one or more logical operators.
Operands can be specified in four ways:
  • Placeholders: OBsmith placeholders for random values (e.g., numberLiteral, booleanLiteral) or random references (numberReference, booleanReference).
  • Identifiers: Direct references to existing variables in scope (e.g., score, isReady).
  • Literals: Any literal, such as a string, number, or boolean.
  • Nested Expressions: An operand can be the result of another expression generator call. This allows us to construct arbitrarily complex and deeply nested program logic from a readable, functional representation. The program generator recursively parses these expressions from the inside out.

### v2

@@ Abstract.
JavaScript obfuscators are widely deployed to protect intellectual property and resist reverse engineering, yet their correctness has been largely overlooked compared to performance and resilience. Existing evaluations typically measure resistance to deobfuscation, leaving the critical question of whether obfuscators preserve program semantics unanswered. Incorrect transformations can silently alter functionality, compromise reliability, and erode security—undermining the very purpose of obfuscation. To address this gap, we present OBsmith, a novel framework to systematically test JavaScript obfuscators using large language models (LLMs). OBsmith leverages LLMs to generate program sketches—abstract templates capturing diverse language constructs, idioms, and corner cases—which are instantiated into executable programs and subjected to obfuscation under different configurations. Besides LLM-powered sketching, OBsmith also employs a second source: automatic extraction of skeletons from real programs. This extraction path enables more focused testing of project-specific features and lets developers inject domain knowledge into the resulting test cases. OBsmith uses two techniques to derive test oracles: (i) reference-oriented equivalence testing, which takes the original program as reference oracle (ground truth) and checks whether the obfuscated version preserves equivalent functionality, and (ii) metamorphic testing, which applies semantics-preserving transformations to the original program and checks if obfuscation violates expected behavior.
We evaluate OBsmith on two widely used obfuscators, Obfuscator.IO and JS-Confuser, generating 600 sketches using six popular LLMs. OBsmith fills these sketches and generates over 3,000 candidate programs and obfuscates them across seven obfuscation configurations. OBsmith uncovers 11 previously unknown correctness bugs. Under an equal program budget, five general purpose state-of-the-art JavaScript fuzzers (FuzzJIT, Jsfunfuzz, Superion, DIE, Fuzzilli) failed to detect these issues, highlighting OBsmith’s complementary focus on obfuscation-induced misbehavior. An ablation shows that all components except our generic MRs contribute to at least one bug class; the negative MR result suggests the need for obfuscator-specific metamorphic relations. Our results also seed a discussion on how to balance obfuscation presets and performance cost. We envision OBsmith as an important step towards automated testing and quality assurance of obfuscators and other semantic-preserving toolchains.
@@
copyright: rightsretaineddoi: 10.1145/3798204journalyear: 2026journal: PACMPLjournalvolume: 10journalnumber: OOPSLA1article: 96publicationmonth: 4ccs: Software and its engineering Formal software verificationccs: Software and its engineering Correctnessccs: Software and its engineering Completenessccs: Software and its engineering Software verificationccs: Computing methodologies Artificial intelligenceccs: General and reference Metricsccs: General and reference Evaluation
@@ 1. Introduction
By design, an effective obfuscator is expected to complicate the original source code without altering its intended functionality. The primary goal is to make the code substantially more difficult for reverse engineers and automated analysis tools to understand, thus safeguarding proprietary or sensitive information. Ensuring functional equivalence between obfuscated and original code is critical, as obfuscation may introduce unintended behaviors, potentially undermining software reliability or security. There is still a lack of research on evaluating obfuscators on pure JavaScript code. Previous research on JavaScript obfuscation largely focuses on evaluating the effectiveness of obfuscation techniques or detecting if code is obfuscated. However, comprehensive testing frameworks that validate both functional correctness and runtime performance impacts of JavaScript obfuscators remain underexplored. Recent research by Skolka et al. (Skolka et al., 2019) offers a comprehensive evaluation of JavaScript minifiers and obfuscators, with particular emphasis on their robustness when applied to client-side scripts. The authors primarily relied on unit tests to assess the effectiveness of various obfuscators. Their findings reveal that these tools often exhibit insufficient resilience. A core challenge of client-side script is that client-side JavaScript is inherently dependent on complex factors such as user interactions, asynchronous events, and integration with real-world browser APIs (e.g., DOM, cookies, localStorage). These dependencies are difficult to replicate or comprehensively test using automated methods. As a result, automated source code transformations—such as minification and obfuscation—may inadvertently introduce bugs in client-side. Moreover, prior studies have omitted the evaluation of some of the most widely adopted obfuscators, such as Obfuscator.IO (obi, 2024) and JS-Confuser (jsc, 2022), as indicated by a recent Google survey (Jiang et al., 2026). This omission constitutes another important research gap, particularly given the pervasive use of these tools in the software industry. To address this limitation, the present study aims to instrument and evaluate pure JavaScript code that is executable in both Node.js and browser environments. Addressing this research gap is crucial to ensuring that JavaScript obfuscators can reliably achieve their intended security and functionality objectives without inadvertently compromising software behavior.
@@
General purpose compiler testing approaches lack a systematically-enhanced test oracle to capture the JavaScript obfuscator bugs. Off-the-shelf compiler fuzzers assume closed-world, deterministic pipelines whose goal is transparent semantics preservation; they rely on crash oracles or n-version agreement (e.g., GCC vs. Clang). JavaScript obfuscators instead preserve while disguising semantics, inject seed-dependent randomness and self-defense checks, and interact with host features that general fuzzers neither generate nor control. Bugs often surface in JS-specific corners—eval function, dynamic code loading, scoping/hoisting, prototype mutations, getters/setters, proxies—combined with obfuscation patterns like control-flow flattening or lazy string decoding. These properties defeat differential comparisons across tools and make crashes a poor proxy for silent miscompilations. Therefore, we introduce reference-oriented equivalence testing, which takes the original program as reference oracle (ground truth) and checks whether the obfuscated version preserves equivalent functionality. Checked correctness includes return values, exceptions, scheduling-sensitive effects, and observable host-API interactions. OBsmith generates JS-centric, event-aware sketches, controls nondeterminism via seed management, executes in sandboxed Node and browser-like environments, and is configuration-aware to sweep obfuscator options. Lightweight instrumentation records variable states and control-flow to localize divergences. This semantics-, host-, and transformation-aware design exposes obfuscation-specific bugs that general compiler fuzzers systematically miss. Obfuscators differ from compilers because they preserve but disguise semantics. They often use tricks (control-flow flattening, dead code) that do not appear in compiler pipelines. Hence, domain-specific sketching/testing is required.
@@
We propose OBsmith, a novel testing framework specifically designed for evaluating JavaScript obfuscators. OBsmith leverages sketching techniques to systematically generate diverse test programs and employs reference-oriented equivalence testing to identify discrepancies introduced by obfuscation processes. Sketching and differential testing methodologies have previously demonstrated effectiveness in compiler testing, enabling the detection of subtle semantic bugs and inconsistencies. Extensive literature on compiler testing underscores the value and efficacy of integrating domain-specific knowledge into test-case generation. Inspired by these insights, we adopt sketching to incorporate JavaScript-specific domain knowledge into our framework, thereby enhancing the capability of OBsmith to detect functional deviations and performance degradations caused by obfuscators.
@@
OBsmith uses LLM to generate sketches and uses Babel (bab, 2014) to implement the reference-oriented equivalence testing mechanism. Trained on extensive textual corpora, LLMs inherently encapsulate rich domain-specific knowledge, making them particularly suitable for generating sketches in automated test-case generation. Leveraging their ability to understand syntactic structures and semantic constraints, LLMs can systematically produce representative code sketches that reflect realistic programming patterns and corner cases. Integrating LLM-based sketch generation into OBsmith frameworks enhances test diversity and coverage with low resource, thus enabling the effective discovery of subtle functional inconsistencies and edge-case defects in software, such as those potentially introduced by JavaScript obfuscators.
@@
A critical component of OBsmith is its integrated program generator, which is engineered to rigorously explore different execution paths and expose the semantic equivalence between the original source program and its obfuscated counterpart. This step is indispensable, as the intricate nature of obfuscation transformations carries an inherent risk of introducing functional regressions or altering program behavior in subtle ways. To mitigate this risk, OBsmith contains a sketch filling algorithm to generate a comprehensive set of inputs to test different execution paths. OBsmith uses a program enhancer to record global and local variables and the program’s control flow. For each input, OBsmith asserts that the original and obfuscated program versions produce identical outputs and exhibit equivalent observable side-effects. This process provides a strong, empirical guarantee that the core logic and functionality of the application remain unaltered post-obfuscation, thereby enhancing the trustworthiness and practical deployability of our approach.
@@
Unlike the traditional differential testing setup used in compiler testing – where multiple variants of a system are run on the same input and their outputs compared for inconsistencies – OBsmith  relies on a reference oracle to serve as ground truth. In particular, we treat the original, unobfuscated program as the reference oracle for correct behavior. This direct comparison against a known-correct reference, termed reference-oriented equivalence testing, allows for a precise and conclusive assessment of the correctness of the obfuscation transformations, ensuring that they have not introduced functional changes.
@@
We evaluate OBsmith on two widely used obfuscators, Obfuscator.IO (obi, 2024) and JS-Confuser (jsc, 2022), generating 600 sketches using six popular LLMs. OBsmith fills these sketches and generates over 3,000 candidate programs and obfuscates them across seven obfuscation configurations. OBsmith uncovers 11 previously unknown correctness bugs. Under an equal program budget, five general purpose state-of-the-art JavaScript fuzzers (FuzzJIT, Jsfunfuzz, Superion, DIE, Fuzzilli) failed to rediscover these issues, highlighting OBsmith’s complementary focus on obfuscation-induced misbehavior. We also discuss obfuscation affect on file size, run time, and memory usage in discussion (§5).
@@
  • Novelty. This paper introduces OBsmith, the first LLM-powered framework for systematically testing JavaScript obfuscators. OBsmith introduces an LLM-driven, sketch-based generator (mixing LLM-created and automatically extracted sketches) plus a PL-aware oracle that combines reference-oriented equivalence testing of outputs/exceptions/termination with lightweight instrumentation and exploratory metamorphic tests.
  • Implementation. We implement OBsmith with 2 sketch sources: (i) leverages multi-agent LLM system to automatically generate diverse sketches and (ii) uses existing JavaScript programs to automatically extract sketches. To solve the oracle problem, OBsmith applies reference-oriented equivalence testing and metamorphic testing to evaluate the correctness of obfuscation.
@@ 2. Sketch Example
In this paper, a sketch refers to a “program with holes,” where key components - such as expressions, variables, or literals - are left as placeholders to be filled in later. Sketches, also known as “skeletal programs” or “templates,” have been shown in previous work to be an effective mechanism for introducing domain knowledge into compiler testing (Zang et al., 2023; Zhang et al., 2017; Wang et al., 2018). By abstracting certain program elements, sketches allow for a flexible yet systematic exploration of diverse code scenarios that are effective in exposing compiler bugs. The left part of Fig. 1 is an example of sketch with holes (e.g. NumberLiteral, NumberReference) and the middle is the corresponding filled code.
@@
JavaScript obfuscators take two inputs: the program under obfuscation and the obfuscation configuration. Once we obfuscate the concrete program with different obfuscators and different configurations, we got the multiple obfuscated programs. Then we conduct a reference-oriented equivalence testing on these programs to compare if they have the same functionality. Unlike traditional differential testing in compiler testing, we have ground truth (unobfuscated program). Here the concrete program serve as ground truth and we compare the running result to find bugs. Furthermore, OBsmith enhances programs by recording the program’s workflow and data flow, which effectively reflects the program functionality and enables us to easily implement reference-oriented equivalence testing. An example of found bug is shown in the right part of Fig. 1. Here the variable x is reported as undefined during execution (ReferenceError: x is not defined). This error highlights a critical functionality bug introduced by JS-Confuser, demonstrating how OBsmith effectively detects discrepancies between original program and obfuscated program.
@@ 3. OBsmith Approach
In this section, we present OBsmith, an automated framework that leverages LLMs to generate sketches and use sketches to test JavaScript obfuscators. OBsmith constructs a multi-step pipeline to test JavaScript obfuscators, its overall workflow is shown in Fig. 2.
@@
Then LLM-generated and extracted sketches are used in program generation3.5) to get concrete JavaScript programs. In program generation stage, OBsmith (i) fills the sketch by filling all placeholders and generate concrete expressions in the sketch; and (ii) enhances the filled program to log the program’s control flow and data flow. The output of program generation stage is a set of enhanced candidate programs ready for obfuscation.
The last step is use candidate programs and call obfuscators to get multiple obfuscated variants. These variants with corresponding ground truth are inputs to the follow-up testing stage. OBsmith uses a reference oracle for reference-oriented equivalence testing (§3.7), complemented by metamorphic testing (§3.8) for robustness. OBsmith uses different JavaScript obfuscators with different obfuscation configurations to obfuscate candidate programs. The original candidate program and all obfuscated versions are executed within a standard JavaScript engine (Node.js). Unlike traditional differential testing in compiler testing which lacks ground truth, OBsmith uses the original program as ground truth. OBsmith collects the execution outputs of all obfuscated programs and compares these outputs with the ground truth. If OBsmith finds any output inconsistencies, it reports the obfuscator as failed. Our correctness criteria and observational equivalence are defined in §3.6.
@@ 3.1. Sketch Definition
This section formally defines the sketch language used by OBsmith. A sketch is a syntactically valid JavaScript program augmented with placeholders that abstract over literals, variable references, and expressions. Sketches are instantiated into concrete programs by a separate sketch-filling procedure described in Section 3.5.
Overview
The sketch language is designed to satisfy two goals: (1) every sketch should be parsable as a JavaScript program by standard tooling (e.g., Babel), and (2) a single sketch should compactly represent a large family of concrete programs. To this end, placeholders are encoded using ordinary identifiers and function calls, while semantic constraints such as scope resolution and operator selection are enforced during sketch filling rather than in the grammar.
EBNF Grammar
Table 1 presents the EBNF grammar of the sketch language. The grammar specifies only the syntactic structure of sketches; it does not encode semantic constraints such as variable scope, type compatibility, or randomized generation strategies.
Table 1. EBNF grammar of sketch expressions embedded in JavaScript programs.
Nonterminal Production
Program $\{\,JavaScriptStatement\,\}$
JavaScriptStatement any JavaScript statement
Expression Atom
$\mid$ ArithmeticExpr
$\mid$ RelationExpr
$\mid$ LogicExpr
Atom Identifier
$\mid$ Literal
$\mid$ Placeholder
Literal NumberLiteral
$\mid$ BooleanLiteral
Placeholder numberLiteral
$\mid$ booleanLiteral
$\mid$ numberReference
$\mid$ booleanReference
ArithmeticExpr arithmetic(Expression, Expression, OperatorList)
RelationExpr relation(Expression, Expression, OperatorList)
LogicExpr logic(Expression, Expression, OperatorList)
OperatorList Operator
$\mid$ Operator , OperatorList
Table 2. Operator categories used in expression placeholders.
Category Operators
ArithmeticOp +, -, *, /, …
RelationalOp <, >, <=, >=, ==, ===, …
LogicalOp &&, ||
Table 3. Lexical categories used in the sketch grammar.
Token Description
Identifier JavaScript identifier
NumberLiteral JavaScript numeric literal
BooleanLiteral true $\mid$ false
OtherJSStatement Any JavaScript statement not involving placeholders
Operator Categories
Operators appearing in expression placeholders are drawn from the categories shown in Table 2. Operator lists specify alternative operators from which one is selected during sketch filling.
Lexical Categories
Table 3 summarizes the lexical categories referenced in the grammar. These categories follow standard JavaScript syntax and are not extended by the sketch language.
Discussion
The EBNF grammar defines the syntactic shape of sketches and ensures that every sketch is a valid JavaScript program. It intentionally omits semantic constraints such as variable scope, operand compatibility, and operator selection. These aspects are handled during sketch filling, where placeholders are resolved and expression generators are instantiated into concrete JavaScript expressions. This separation allows the sketch language to remain simple while enabling flexible and diverse program generation.